4xsquared+25ysquared-100=0 put in standard form and graphed
Answer by fractalier(6550)

From
4x^2 + 25y^2 - 100 = 0
add 100 and then divide by 100
4x^2 + 25y^2 = 100
x^2 / 25 + y^2 / 4 = 1
This is an ellipse with semi-major axis of 5 and semi-minor axis of 2.
Intercepts at (5, 0), (-5, 0), (0, 2) and (0, -2).